OpenStudy (anonymous):

Which of the following is the slope between the points (-1, -5) and (7, -3)? 4 one fourth -negative one fourth -4

OpenStudy (anonymous):

\[ m = y _{2} - y _{1} \div x _{2} - x _{1}\]

OpenStudy (anonymous):

-3 - (-5) = -3 + 5 = 2 7 - (-1) = 7 +1 = 8 2/8 = 1/4